OpenStudy (anonymous):

2. A company has fixed operating costs of $1,840.00 per month with a production cost of $12.32 per unit. If each unit brings $25.01 in revenue, which of the following equations represents the profit for the month? (Let x represent the number of units made per month and y represent the total profit for the month.) y = 12.69x - 1,840 y = 25.01x - 1,840 y = 12.32x - 1,840 y = 37.33x - 1,840

OpenStudy (hoblos):

the profit is the difference between the money gained and the money paid total money gained = 25.01x total money paid = 1,840 + 12.32x so y= 25.01x -(1,840 + 12.32x) = 25.01x -1,840 - 12.32x) = 12.69x - 1,840

OpenStudy (anonymous):

thank you (:

OpenStudy (hoblos):

glad to help :)